Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

bag of baloon: probability!!!

Expert replies
by acorra » Sat Apr 25, 2009 5:52 am
In a bag of baloons, the ratio of red baloons to other colours is 1:5, if 3 baloons are selected from the bag at random, what is the probability that:
a) at least one baloon is red
b) exactly one baloon is red

OA will follow soon
Join the discussion
Source: — Problem Solving |

Re: bag of baloon: probability!!!

by Ian Stewart » Sat Apr 25, 2009 6:59 am
acorra wrote:In a bag of baloons, the ratio of red baloons to other colours is 1:5, if 3 baloons are selected from the bag at random, what is the probability that:
a) at least one baloon is red
b) exactly one baloon is red

OA will follow soon
You cannot answer this kind of question if you only know the ratio of red balloons to other balloons. Take question a), for example. Notice that we either get zero red balloons, or we get at least one red balloon, so

Prob(0 red balloons) + Prob(1 or more red balloons) = 1

Prob(1 or more red balloons) = 1 - Prob(0 red balloons)

So, if we can find the probability that all three balloons are not red, we can find the probability that at least one balloon is red by subtracting from one.

Now, say we have exactly 1 red balloon, and exactly 5 non-red balloons. The probability that the first is not red is 5/6, that the second is again not red is 4/5, and that the third is again not red is 3/4. So the probability that all three are non-red is

(5/6)(4/5)(3/4) = 1/2

So the probability of 'at least one red' is 1 - 1/2 = 1/2

On the other hand, if we have 2 red balloons, and 10 non-red balloons, the probability that all are non-red is

(10/12)(9/11)(8/10) = 6/11

and the probability that at least one is red is 5/11.

So the ratio alone is not sufficient to get a numerical answer here - the answer is different depending on the number of balloons in the bag. Of course we could solve the problem using an unknown. If we have x red balloons, we have 5x non-red balloons. The probability of selecting three non-red balloons is therefore:

(5x/6x)[(5x - 1)/(6x - 1)][(5x - 2)/(6x - 2)]

and subtracting that from 1 (and canceling the common factor of x in the numerator and denominator), you could get an algebraic expression for the probability in question, in terms of x, the number of red balloons. But that's about the best you can do here.

Throughout, I've assumed that the selections are made 'without replacement', something which is strongly suggested by the wording of the question. If the selections are made 'with replacement' (i.e. after selecting each balloon, we put it back in the bag before making our next selection), then the answer would be 1 - (5^3/6^3), since the probability would be 5/6 on each selection that a balloon is non-red. As the question is worded, however, that interpretation (that the selections are made with replacement) doesn't make much sense.
For online GMAT math tutoring, or to buy my higher-level Quant books and problem sets, contact me at ianstewartgmat at gmail.com

ianstewartgmat.com
Join the discussion